OpenStudy (anonymous):

x to the 4 minus 8x squared plus 7=0

OpenStudy (campbell_st):

\[x^4 -8x^2 + 7 = 0\] this factorises to \[(x^2 - 7)(x^2 - 1)=0\] which gives \[x = \pm \sqrt{7}... and... x = \pm 1\]

OpenStudy (cwrw238):

x^4 - 8x ^2 + 7 = 0 let x^2 = X X^2 - 8X + 7 = 0 (X - 7)(X - 1) = 0 X = 7 or 1 can you continue with this?